# On the number of linear uniform hypergraphs with girth constraint

### v1

@@ 2 Preliminaries
We adopt the standard asymptotic Landau notation and expressions from [7] throughout. All asymptotics in this paper are with respect to $n\to\infty$. For two positive-valued functions $a(n)$, $b(n)$ on the variable $n$, write $a(n)=O(b(n))$ if there exists a constant $C>0$ such that for all $n$, $|a(n)|\leqslant C\cdot|b(n)|$. Write $a(n)=o(b(n))$ or $a(n)\ll b(n)$ if eventually $b(n)>0$ and $\lim_{n\rightarrow\infty}a(n)/b(n)$ exists and equals 0; $a(n)=b(n)(1+o(1))$ or $a(n)\sim b(n)$ means $\lim_{n\rightarrow\infty}a(n)/b(n)=1$. Denote $a(n)=\Omega(b(n))$ if eventually $a(n)>0$ and $b(n)=O(a(n))$. Further, $a(n)=\Theta(b(n))$ if both $a(n)=O(b(n))$ and $b(n)=\Omega(a(n))$. Let $a=b\pm c$ be short for $a\in[b-c,b+c]$. Equations containing $\pm$ are always to be interpreted from left to right, i.e. $b_{1}\pm c_{1}=b_{2}\pm c_{2}$ means that $b_{1}-c_{1}\geqslant b_{2}-c_{2}$ and $b_{1}+c_{1}\leqslant b_{2}+c_{2}$. Moreover, given a set S{S} and integers $a,b\geqslant 0$, we write ${S}\choose b$} for the collection of all $b$-sets of S{S}; in particular, ${S}\choose b$$\emptyset$ and ${a\choose b}=0$ if b>|S|b>|{S}| and $b>a$, $a\wedge b$ denotes the minimum of $a$ and $b$. All logarithms are natural. The floor and ceiling signs are omitted whenever they are not crucial.
@@
Rescale the number of steps $i$ to be $t=t_{i}=\frac{i}{n^{2}}$ such that $t_{i+1}=t+\frac{1}{n^{2}}$ and every function of the discrete variable $i$ can be viewed as a continuous function of $t$. It is sometimes more convenient to think of our random variables as depending on a continuous variable $t$ rather than the discrete variable $i$. We pass between these interpretations without comments. The main tool in the proof of our main results is the differential equations method, which applies a pseudo-random heuristic to divine the trajectories of a group of graph parameters that govern the evolution of the process. It plays a central role in the understanding of several other constrained random processes to produce interesting combinatorial objects [4, 5, 6, 7, 16, 18, 20, 21, 23]. The key in applying the method is to find a collection of random variables containing |X(i)||{X}(i)| whose one step change can be expressed in terms of other variables in the collection. These expressions form a system of difference equations and the solution to the initial value problem for the corresponding system of differential equations gives a natural guess for the expected trajectories of the variables. The one-step change and boundedness hypotheses for each tracked variable are analyzed, then the last step is to apply a martingale concentration inequality and a union bound to prove that all of the variables in the collection are concentrated around their expected trajectories.
@@
where the second equality is derived by applying Taylor’s theorem with remainder in Lagrange form on ${x}(t+\frac{1}{n^{2}})-{x}(t)$ and $\epsilon_{x}(t+\frac{1}{n^{2}})-\epsilon_{x}(t)$ until the second order. Based on the estimated expression $x(t)$ and error function $\epsilon_{x}(t)$ into the equations (2.1) and (2.2), we will show that the terms $\mathbb{E}[\Delta{X}(i)\,|\mathcal{F}_{i}]$ and $\frac{{x}{{}^{\prime}}(t)}{n^{2}}$ cancel each other out; $\frac{\epsilon_{x}^{\prime}(t)}{n^{2}}$ dominates the remaining terms such that $\mathbb{E}[\Delta{X}^{\pm}(i)\,|{\mathcal{F}}_{i}]\leqslant 0$ and $\mathbb{E}[|\Delta{X}^{\pm}(i)|\,|{\mathcal{F}}_{i}]=O(\frac{|\epsilon_{x}^{\prime}(t)|}{n^{2}})$; then we also have $|\Delta{X}^{\pm}(i)|\leqslant|\Delta{X}(i)|+O(\max\{\frac{|{x}^{\prime}(t)|}{n^{2}},\frac{|\epsilon_{x}^{\prime}(t)|}{n^{2}}\})$ and the initial value ${X}^{\pm}(0)\leqslant-\frac{1}{2}\epsilon_{x}(0)$. Applying Freedman’s martingale concentration inequality [12] (see also [24]) and a union bound of the events ${X}^{\pm}(i)\geqslant 0$ for all tracked variables at every step $i\leqslant M$, we prove that the designated variable ${X}^{\pm}(i)\geqslant 0$ from some $i\leqslant{M}$ has extremely low probability. For this to work, it is essential to strictly characterize $x(t)$ and $\epsilon_{x}(t)$ for each tracked variable, where $\epsilon_{x}(t)$ has a large enough growth rate throughout the process, but $\epsilon_{x}(t)$ must not grow too fast, otherwise we would lose control of X(i){X}(i).
3 High girth $r$-clique removal process
@@ 3.1 Definitions and Variables
It is observed that the structure of Type-$i$ configuration is a linear cycle $C_{i}^{r}$, because it is linear with girth exactly $\ell$, which implies that ${v}_{L}=(r-1)\cdot{e}_{L}$ for any $L\in\mathcal{L}$, and ${v}_{G}\geqslant(r-1)\cdot{e}_{G}+1$ for any $G\subsetneq L$ with $L\in\mathcal{L}$. For any fixed integer $\ell$, $|\mathcal{L}|=O(1)$.
The random greedy high girth $r$-clique removal process starts with a complete graph on vertex set $[n]$, denoted by $\mathbb{G}(0)$, and $\mathbb{G}(i+1)$ is the remaining graph from $\mathbb{G}(i)$ by selecting one $r$-clique uniformly at random out of all available $r$-cliques in $\mathbb{G}(i)$ and deleting all its edges from the edge set E(i){E}(i) of $\mathbb{G}(i)$, where an $r$-clique is called available in $\mathbb{G}(i)$ means that adding the vertex set of this $r$-clique as an edge does not produce an $r$-graph in $\mathcal{L}$ with the edges generated by the vertices of previously chosen available $r$-cliques. Let ${M}=\min\{i:\mathbb{G}(i){\text{ is}\textit{ available }r{\text{-}\textit{clique}\text{ free}}}\}$. This process is equivalently viewed as creating a random linear $r$-graph with girth greater than $\ell$, that is, the random greedy high girth linear $r$-uniform hypergraph process. Beginning with the empty $r$-graph $\mathbb{H}(0)$ on the same vertex set $[n]$ we sequentially set $\mathbb{H}({i+1})=\mathbb{H}(i)+\boldsymbol{e}_{i+1}$, where the added $r$-set $\boldsymbol{e}_{i+1}$ is the vertex set of the chosen available $r$-clique in $\mathbb{G}(i)$, and $\boldsymbol{e}_{i+1}$ is also called to be an available $r$-set or available edge for $\mathbb{H}(i)$. The process terminates at a maximal linear $r$-graph $\mathbb{H}({M})$ with girth larger than $\ell$ and no available $r$-sets or available edges. At time $i$, graphs are defined with respect to edge sets on the vertex set $[n]$, where both ${\mathbb{G}}(i)$ and $\mathbb{H}(i)$ depend on the outcomes of the process up to this point, and ${\mathbb{G}}(i)$ is the graph given by the pairs of vertices that do not appear in any edge of $\mathbb{H}(i)$. It will be convenient to study the relation between $\mathbb{G}(i)$ and $\mathbb{H}(i)$ via
@@
For $2\leqslant m\leqslant r$, ${u}\in[n]$ and $\boldsymbol{f}_{m}=\{{u}_{1},\cdots,{u}_{m}\}\in\binom{[n]}{m}$, let ${N}_{{u}}(i)=\{{x}\in[n]:\{x,u\}\in{E}(i)\}$, ${N}_{\boldsymbol{f}_{m}}(i)=\cap_{j=1}^{m}{N}_{{u}_{j}}$ and ${K}_{m}(i)$ be the set of vertex sets of $m$-cliques in ${\mathbb{G}}(i)$. Due to the nature of the random greedy high girth $r$-clique removal process, it should come as no surprise that the most important variable for us to track is the number of available $r$-cliques in $\mathbb{G}(i)$. To this end, define Q(i){Q}(i) to be the set of vertex sets of available $r$-cliques in ${\mathbb{G}}(i)$. Our goal is to estimate the random variable |Q(i)||{Q}(i)|. Note that $|{Q}(0)|=\binom{n}{r}$, and at termination of the process |Q(M)|=0|{Q}({M})|=0. Fix one $\boldsymbol{f}_{m}\in{K}_{m}(i)$, and define the set ${Y}_{\boldsymbol{f}_{m}}(i)$ to be
@@ 3.2 Estimates on the variables
A pseudo-random heuristic for divining the evolution of variables plays a central role in the understanding of graph processes. The following idea of estimating on the key variables has been applied in the environments of some interesting combinatorial objects [4, 5, 6, 7, 16, 18]. We extend the ideas in [16, 7] to show the expected trajectories for each tracked variable of the high girth linear uniform hypergraph process with a higher uniformity.
Suppose that the events that each available $r$-set, denoted by $\boldsymbol{f}$, is chosen to become an edge in $\mathbb{H}(i)$ are independent with probability $\pi_{i}=\pi(t)$ without considering the forbidden configurations in $\mathcal{L}$ defined in Definition 3.1, that is, $\mathbb{P}[\boldsymbol{f}\in\mathbb{H}(i)]=\pi_{i}$, which implies that $\mathbb{H}(i)$ is regarded as the binomial random $r$-graph $\mathcal{H}_{r}(n,\pi_{i})$. Recalling that $t=t_{i}=\frac{i}{n^{2}}$ and in view of $|\mathbb{H}(i)|=i\sim\binom{n}{r}\cdot r!t/n^{r-2}$, we have
@@
From another viewpoint, it means that the graph ${\mathbb{G}}(i)$ is the random graph whose variables are roughly the same as they are in the binomial random graph $\mathcal{G}(n,p_{i})$ with
@@
Given any $\mu>0$ and fixed integers $r,\ell\geqslant 3$, there exist $\lambda>0$ and $\alpha\in(0,1)$ such that, for any ${\boldsymbol{f}}_{m}\in{K}_{m}(i)$ with $2\leqslant m\leqslant r-1$, ${\boldsymbol{f}}\in{Q}(i)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$, with probability at least $1-n^{-\mu}$,
@@
Theorem 3.4 is proved in Section 4 and we always tacitly assume $0\leqslant i\leqslant{M}$ with M{M} defined in (3.18). It implies that the process produces a linear $r$-graph of size at least M{M} with its girth greater than $\ell$. Theorem 3.4 verifies Theorem 1.2 with room to spare in the power of the logarithmic factor. We make no attempt to optimize the constants $\lambda$, $\alpha$ and the coefficients $\beta_{k}$ for any $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$. There are many choices of them that can be balanced to satisfy certain inequalities. For example, for given positive integers $r$, $\ell$ and $k$, we choose these constants to satisfy the equations
@@
where $\alpha_{0}=\frac{\ell-2}{\ell-1}$, $p_{{M}}$ and $\xi_{{M}}$ are constants when $i=M$ defined in (3.5) and (3.9). We will see these choices are sufficient for our proof of Theorem 3.4 in next section. We do not replace them with their actual values, which is for the interest of understanding the role of these constants played in the calculations.
@@
According to (3.5) and (3.9), note that $p_{i},\xi_{i}=\Theta(1)$ when $t\in[0,t_{{M}}]$ with $t_{{M}}$ defined in (3.26). Unfortunately, it is not allowed to choose $t_{M}=n^{-1+\frac{1}{\ell-1}}$ in the proof of our main results in Section 4 because we can not remove $\sigma_{i}$ or change the power of $\sigma_{i}$ in (3.22)-(3.24). Based on the analysis and calculations in Section 4, the factor $\sigma_{i}$ is necessary to show the sequences of auxiliary random variables for each tracked variable are supermartingales. However, $\xi_{i}\rightarrow 0$ as $t=n^{-1+\frac{1}{\ell-1}+\varepsilon}$ for any arbitrarily small $\varepsilon>0$, which implies it is impossible to apply this approach to show a better lower bound of $\textmd{ex}_{L}(n,r,\ell)$ than $\Omega(n^{1+\frac{1}{\ell-1}})$.
@@
4 Analysis of the high-girth process
For any integer $0\leqslant i<{M}$, let $\mathcal{G}(i)$ denote the event that the estimates in Theorem 3.4 hold for any $\boldsymbol{f}_{m}\in{K}_{m}(j)$ with $2\leqslant m\leqslant r-1$, $\boldsymbol{f}\in{Q}(j)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$ when $0\leqslant j\leqslant i$. The initial conditions in $\mathbb{G}(0)$ satisfy the equations in (3.15)-(3.17), which implies that the event $\mathcal{G}(0)$ always holds. Assume that the event $\mathcal{G}(i)$ holds with probability at least $1-n^{-\mu}$ for some $\mu>0$. Following the approach in [16, 7] to consider sparse Steiner triple systems, we analyze the random greedy high girth linear $r$-uniform hypergraph process in this section and finally show that the event $\mathcal{G}({{M}})$ holds with high probability to complete the proof of Theorem 3.4.

### v2

@@ 2 Preliminaries
We adopt the standard asymptotic Landau notation and expressions from [8] throughout. All asymptotics in this paper are with respect to $n\to\infty$. For two positive-valued functions $a(n)$, $b(n)$ on the variable $n$, write $a(n)=O(b(n))$ if there exists a constant $C>0$ such that for all $n$, $|a(n)|\leqslant C\cdot|b(n)|$. Write $a(n)=o(b(n))$ or $a(n)\ll b(n)$ if eventually $b(n)>0$ and $\lim_{n\rightarrow\infty}a(n)/b(n)$ exists and equals 0; $a(n)=\omega(b(n))$ if eventually $a(n)>0$ and $b(n)=o(a(n))$; $a(n)=b(n)(1+o(1))$ or $a(n)\sim b(n)$ means $\lim_{n\rightarrow\infty}a(n)/b(n)=1$. Denote $a(n)=\Omega(b(n))$ if $b(n)=O(a(n))$. Further, $a(n)=\Theta(b(n))$ if both $a(n)=O(b(n))$ and $b(n)=\Omega(a(n))$. Let $a=b\pm c$ be short for $a\in[b-c,b+c]$. Equations containing $\pm$ are always to be interpreted from left to right, i.e. $b_{1}\pm c_{1}=b_{2}\pm c_{2}$ means that $b_{1}-c_{1}\geqslant b_{2}-c_{2}$ and $b_{1}+c_{1}\leqslant b_{2}+c_{2}$. Moreover, given a set S{S} and integers $a,b\geqslant 0$, we write ${S}\choose b$} for the collection of all $b$-sets of S{S}; in particular, ${S}\choose b$$\emptyset$ and ${a\choose b}=0$ if b>|S|b>|{S}| and $b>a$; $a\wedge b$ denotes the minimum of $a$ and $b$. All logarithms are natural. The floor and ceiling signs are omitted whenever they are not crucial.
@@
Rescale the number of steps $i$ to be $t=t_{i}=\frac{i}{n^{2}}$ such that $t_{i+1}=t+\frac{1}{n^{2}}$ and every function of the discrete variable $i$ can be viewed as a continuous function of $t$. It is sometimes more convenient to think of our random variables as depending on a continuous variable $t$ rather than the discrete variable $i$. We pass between these interpretations without comments. The main tool in the proof of our main results is the differential equations method, which applies a pseudo-random heuristic to divine the trajectories of a group of graph parameters that govern the evolution of the process. It plays a central role in the understanding of several other constrained random processes to produce interesting combinatorial objects [4, 5, 7, 8, 19, 20, 23, 25, 28, 29, 31]. The key in applying the method is to find a collection of random variables containing |X(i)||{X}(i)| whose one step change can be expressed in terms of other variables in the collection. These expressions form a system of difference equations and the solution to the initial value problem for the corresponding system of differential equations gives a natural guess for the expected trajectories of the variables. The one-step change and boundedness hypotheses for each tracked variable are analyzed, then the last step is to apply a martingale concentration inequality and a union bound to prove that all of the variables in the collection are concentrated around their expected trajectories.
@@
where the second equality is derived by applying Taylor’s theorem with remainder in Lagrange form on ${x}(t+\frac{1}{n^{2}})-{x}(t)$ and $\epsilon_{x}(t+\frac{1}{n^{2}})-\epsilon_{x}(t)$ until the second order. Based on the estimated expression $x(t)$ and error function $\epsilon_{x}(t)$ into the equations (2.1) and (2.2), we will show that the terms $\mathbb{E}[\Delta{X}(i)\,|\mathcal{F}_{i}]$ and $\frac{{x}{{}^{\prime}}(t)}{n^{2}}$ cancel each other out; $\frac{\epsilon_{x}^{\prime}(t)}{n^{2}}$ dominates the remaining terms such that $\mathbb{E}[\Delta{X}^{\pm}(i)\,|{\mathcal{F}}_{i}]\leqslant 0$ and $\mathbb{E}[|\Delta{X}^{\pm}(i)|\,|{\mathcal{F}}_{i}]=O(\frac{|\epsilon_{x}^{\prime}(t)|}{n^{2}})$. We also show that $|\Delta{X}^{\pm}(i)|\leqslant|\Delta{X}(i)|+O(\max\{\frac{|{x}^{\prime}(t)|}{n^{2}},\frac{|\epsilon_{x}^{\prime}(t)|}{n^{2}}\})$ and the initial value ${X}^{\pm}(0)\leqslant-\frac{1}{2}\epsilon_{x}(0)$. Applying Freedman’s martingale concentration inequality [15] (see also [32]) and a union bound of the events ${X}^{\pm}(i)\geqslant 0$ for all tracked variables at every step $i\leqslant M$, we prove that the designated variable ${X}^{\pm}(i)\geqslant 0$ from some $i\leqslant{M}$ has extremely low probability. For this to work, it is essential to strictly characterize $x(t)$ and $\epsilon_{x}(t)$ for each tracked variable, where $\epsilon_{x}(t)$ has a large enough growth rate throughout the process, but $\epsilon_{x}(t)$ must not grow too fast, otherwise we would lose control of X(i){X}(i).
3 High linear-girth $r$-clique removal process
@@ 3.1 Definitions and Variables
It is observed that the structure of Type-$i$ configuration is a linear cycle $C_{i}^{r}$, because it is linear with linear girth exactly $\ell$, which implies that ${v}_{L}=(r-1)\cdot{e}_{L}$ for any $L\in\mathcal{L}$, and ${v}_{G}\geqslant(r-1)\cdot{e}_{G}+1$ for any $G\subsetneq L$ with $L\in\mathcal{L}$. For any fixed integer $\ell$, we have $|\mathcal{L}|=O(1)$. The definition here is helpful for analysing the probability of availability for each vertex set of an $r$-clique in (3.8), and the ideal expected values of these tracked variables in (3.10), (3.13) and (3.14). The form in this definition also facilitates the analysis of two groups of auxiliary enumerative parameters in Section 5.
The random greedy high linear-girth $r$-clique removal process starts with a complete graph on vertex set $[n]$, denoted by $\mathbb{G}(0)$, and $\mathbb{G}(i+1)$ is the remaining graph from $\mathbb{G}(i)$ by selecting one $r$-clique uniformly at random out of all available $r$-cliques in $\mathbb{G}(i)$ and deleting all its edges from the edge set E(i){E}(i) of $\mathbb{G}(i)$, where an available $r$-clique in $\mathbb{G}(i)$ means that adding the vertex set of this $r$-clique as an edge does not produce an $r$-graph in $\mathcal{L}$ with the edges generated by the vertices of previously chosen available $r$-cliques. Let ${M}=\min\{i:\mathbb{G}(i){\text{ is}\textit{ available }r{\text{-}\textit{clique}\text{ free}}}\}$. This process is equivalently viewed as creating a random linear $r$-graph with linear girth greater than $\ell$, that is, the random greedy high linear-girth linear $r$-uniform hypergraph process. Beginning with the empty $r$-graph $\mathbb{H}(0)$ on the same vertex set $[n]$ we sequentially set $\mathbb{H}({i+1})=\mathbb{H}(i)+\boldsymbol{e}_{i+1}$, where the added $r$-set $\boldsymbol{e}_{i+1}$ is the vertex set of the chosen available $r$-clique in $\mathbb{G}(i)$, and $\boldsymbol{e}_{i+1}$ is also called to be an available $r$-set or available edge for $\mathbb{H}(i)$. The process terminates at a maximal linear $r$-graph $\mathbb{H}({M})$ with linear girth greater than $\ell$ and no available $r$-sets or available edges. At time $i$, graphs are defined with respect to edge sets on the vertex set $[n]$, where both ${\mathbb{G}}(i)$ and $\mathbb{H}(i)$ depend on the outcomes of the process up to this point, and ${\mathbb{G}}(i)$ is the graph given by the pairs of vertices that do not appear in any edge of $\mathbb{H}(i)$. It will be convenient to study the relation between $\mathbb{G}(i)$ and $\mathbb{H}(i)$ via
@@
For $2\leqslant m\leqslant r$, ${u}\in[n]$ and $\boldsymbol{f}_{m}=\{{u}_{1},\cdots,{u}_{m}\}\in\binom{[n]}{m}$, let ${N}_{{u}}(i)=\{{x}\in[n]:\{x,u\}\in{E}(i)\}$, ${N}_{\boldsymbol{f}_{m}}(i)=\cap_{j=1}^{m}{N}_{{u}_{j}}(i)$ and ${K}_{m}(i)$ be the set of vertex sets of $m$-cliques in ${\mathbb{G}}(i)$. Due to the nature of the random greedy high linear-girth $r$-clique removal process, it should come as no surprise that the most important variable for us to track is the number of available $r$-cliques in $\mathbb{G}(i)$. To this end, define Q(i){Q}(i) to be the set of vertex sets of available $r$-cliques in ${\mathbb{G}}(i)$. Our goal is to estimate the random variable |Q(i)||{Q}(i)|. Note that $|{Q}(0)|=\binom{n}{r}$, and at termination of the process |Q(M)|=0|{Q}({M})|=0. Fix one $\boldsymbol{f}_{m}\in{K}_{m}(i)$, and define the set ${Y}_{\boldsymbol{f}_{m}}(i)$ to be
@@ 3.2 Estimates on the variables
A pseudo-random heuristic for divining the evolution of variables plays a central role in the understanding of graph processes. The following idea of estimating on the key variables has been applied in the environments of some interesting combinatorial objects [4, 5, 7, 8, 19, 23, 25, 31]. We extend the ideas in [19, 8] to show the expected trajectories for each tracked variable of the high linear-girth linear uniform hypergraph process in a higher uniformity.
Suppose that the events that each available $r$-set, denoted by $\boldsymbol{f}$, is chosen to become an edge in $\mathbb{H}(i)$ are independent with probability $\pi_{i}=\pi(t)$ without considering the forbidden configurations in $\mathcal{L}$ defined in Definition 3.1, that is, $\mathbb{P}[\boldsymbol{f}\in\mathbb{H}(i)]=\pi_{i}$. It implies that $\mathbb{H}(i)$ is regarded as the binomial random $r$-graph $\mathcal{H}_{r}(n,\pi_{i})$, where every $r$-set appears independently with probability $\pi_{i}$. Recalling that $t=t_{i}=\frac{i}{n^{2}}$ and in view of $|\mathbb{H}(i)|=i\sim\binom{n}{r}\cdot r!t/n^{r-2}$, we have
@@
From another viewpoint, it means that the graph ${\mathbb{G}}(i)$ is the random graph whose variables are roughly the same as they are in the binomial random graph $\mathcal{G}(n,p_{i})$, where every edge appears independently with probability $p_{i}$ in $\mathcal{G}(n,p_{i})$. Hence, we have
@@
Given any fixed $\mu>0$ and integers $r,\ell\geqslant 3$, there exist $\lambda>0$ and $\alpha\in(0,1)$ such that, for any ${\boldsymbol{f}}_{m}\in{K}_{m}(i)$ with $2\leqslant m\leqslant r-1$, ${\boldsymbol{f}}\in{Q}(i)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$, with probability at least $1-n^{-\mu}$,
@@
Theorem 3.4 is proved in Section 4 and we always tacitly assume $0\leqslant i\leqslant{M}$ with M{M} defined in (3.18). It implies that the process produces a linear $r$-graph of size at least M{M} with its linear girth greater than $\ell$. Theorem 3.4 verifies Theorem 1.2 with room to spare in the power of the logarithmic factor. We make no attempts to optimize the constants $\lambda$, $\alpha$ and the coefficients $\beta_{k}$ for any $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$. There are many choices of them that can be balanced to satisfy certain inequalities. For example, for given positive integers $r$, $\ell$ and $k$, we choose these constants to satisfy the equations
@@
where $\alpha_{0}=\frac{\ell-2}{\ell-1}$, $p_{{M}}$ and $\xi_{{M}}$ are constants when $i=M$ defined in (3.5) and (3.9). We will see that these choices are sufficient for our proof of Theorem 3.4 in next section. We do not replace them with their actual values, which is for the interest of understanding the role of these constants played in the calculations.
@@
According to (3.5) and (3.9), note that the decreasing functions $p_{i}$ and $\xi_{i}$ satisfy $p_{i},\xi_{i}=\Theta(1)$ when $t\in[0,t_{{M}}]$ with $t_{{M}}$ defined in (3.26). Unfortunately, it is not allowed to choose $t_{M}=n^{-1+\frac{1}{\ell-1}}$ in the proof of our main results in Section 4 because we can not remove $\sigma_{i}$ or change the power of $\sigma_{i}$ in (3.22)-(3.24), where the factor $\sigma_{i}$ and its power in (3.23) and (3.24) are sufficient to show the sequence of each tracked variable is a supermartingales based on the analysis and calculations in Section 4.2.1 and 4.2.2, and the factor $\sigma_{i}$ is sufficient to apply Lemma 4.3.1 in Section 4.3.
Remark 3.6.
It is impossible to apply the random greedy analysis to show a better lower bound for the step $i$ of the process than $\Omega(n^{1+\frac{1}{\ell-1}})$ for any pair of fixed integers $r,\ell\geqslant 3$. In fact, $\xi_{i}\rightarrow 0$ as $t=n^{-1+\frac{1}{\ell-1}+\varepsilon}$ for any arbitrarily small $\varepsilon>0$. Hence, $\mathbb{E}[|{Q}(i)|]$ in (3.10), $\mathbb{E}[|{Y}_{\boldsymbol{f}_{m}}(i)|]$ in (3.13) and $\mathbb{E}[|{W}_{\boldsymbol{f},L,k}(i)|]$ in (3.14) all tend to zero such that the analysis here does not work.
@@
4 Analysis of the high linear-girth process
For any integer $0\leqslant i<{M}$, let $\mathcal{G}(i)$ denote the event that the estimates in Theorem 3.4 hold for any $\boldsymbol{f}_{m}\in{K}_{m}(j)$ with $2\leqslant m\leqslant r-1$, $\boldsymbol{f}\in{Q}(j)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$ when $0\leqslant j\leqslant i$. The initial conditions in $\mathbb{G}(0)$ satisfy the equations in (3.15)-(3.17), which implies that the event $\mathcal{G}(0)$ always holds. For any fixed constant $\mu>0$, assume that the event $\mathcal{G}(i)$ holds with probability at least $1-n^{-\mu}$. Following the approach in [19, 8] to consider sparse Steiner triple systems, we analyze the random greedy high linear-girth linear $r$-uniform hypergraph process in this section and finally show that the event $\mathcal{G}({{M}})$ holds with high probability to complete the proof of Theorem 3.4.